OpenStudy (anonymous):

1 1/7 in hours and minutes

OpenStudy (anonymous):

Assuming the 1 and 1/7 is in hours: That is saying you have 1 hour and one seventh of an hour. What's a seventh of an hour? 1/7 * hour But an hour = 60 minutes, so we can say 1/7 * 60 minutes Using a calculator, that's approx = 8.5714 So, 1 1/7 hours approx = 1 hour and 8.5714 minutes.

OpenStudy (anonymous):

60/7 = 8.57 minutes